# Total Failure of Approachability at Successors of Singulars of Countable Cofinality

### v1

@@ 6. The Main Iteration
In this section we will prove the main theorem. Let us first explain the idea behind the construction: The poset $\mathbb{A}(\mu,\kappa)$ forces that for every singular cardinal $\delta>\kappa$ of countable cofinality, there are stationarily many points in $\delta^{+}$ of cofinality $\mu$ which are not approachable. Given this poset, it makes sense to iterate the poset $\mathbb{A}$ for an increasing sequence of supercompact cardinals. The main problem, of course, is showing that the stationarity of the set of non-approachable points of cofinality $\mu$ is preserved. The solution to this problem consists of the following two ideas:
@@
    By Theorem 5.3, after forcing with $\mathbb{A}(\mu,\kappa)$ the stationarity of the set of non-approachable points of cofinality $\mu$ in $\delta^{*}:=(\sup_{n}\delta_{n})^{+}$ is witnessed by models for which we have $\mathbb{P}$-master conditions whenever $\mathbb{P}$ is a $\kappa-$directed closed forcing which preserves $\delta^{*}$, every $\delta_{n}$ and forces $\mathsf{LIP}(\check{\mu},\check{\mu}^{+},\check{\delta}_{n})$ for every $n\in\omega$.
@@
So the main obstacle is showing that $(\mathbb{P},\leq_{0})$ forces many instances of $\mathsf{LIP}$. However, in general working with the direct extension ordering on a Magidor iteration is complicated, so we work with the product of the termspace forcings instead (which suffices by the projection). To show that this forces $\mathsf{LIP}$, we need that our cardinal $\kappa$ is indestructably supercompact. By work of Hamkins and Shelah (see [HS98]), the indestructibility of a supercompact cardinal is destroyed by any nontrivial small forcing. Because of this, we follow every $\mathbb{A}$-iterand with the poset making the next supercompact cardinal indestructible. The following lemma easily follows from well-known work of Laver (see [Lav78]) together with our iteration lemma for strong directed closure (see Lemma 1.23):
@@
We will verify in a series of lemmas that we obtain the desired model after forcing with $\mathbb{P}$. We have to show the following:
@@
In particular, for any $\alpha<\beta$, forcing with $\dot{\mathbb{P}}_{\alpha,\beta}^{G}$ (where $G$ is $\mathbb{P}_{\alpha}$-generic) does not add any new bounded subsets of $\kappa_{\alpha}$ or $\kappa_{\alpha}^{+}$, so the construction stabilizes and the class forcing extension satisfies $\mathsf{ZFC}$. Abusing language, we will refer to this as a forcing extension “by $\mathbb{P}$”, ignoring that $\mathbb{P}$ is a class and not a set.
@@
So we have shown that all $\kappa_{\alpha}$ and $\kappa_{\alpha}^{*}$ remain cardinals. Clearly, whenever $\mu$ does not have this form, $\mu$ is collapsed. So the above set contains precisely all cardinals in the forcing extension. ∎
@@
We want to show that $\mathbb{P}$ gives us exactly the model we need. The proof idea is as follows: The regular cardinals of the forcing extension are precisely the $\kappa_{\alpha}$ for $\alpha$ a successor and $\kappa_{\alpha}^{+}$ for $\alpha$ a limit. For any such cardinal, $\mathbb{A}(\kappa_{\alpha}^{*},\kappa_{\alpha+1})$ forces the existence of stationarily many non-approachable points of cofinality $\kappa_{\alpha}^{*}$ in any successor of a singular cardinal $\kappa_{\alpha+1}$ of countable cofinality. Then all that is left is to show that the tail of the iteration does not destroy that stationarity (since it cannot make points of cofinality $\kappa_{\alpha}^{*}$ approachable due to its closure). For this we use that we have used a guessing function and thus added many partially generic filters to many strongly $\kappa_{\alpha}-$directed closed forcings (namely, those forcings which force enough instances of $\mathsf{LIP}$). We now show that the direct extension ordering on the tail of the iteration can be projected onto from one of those suitable forcings.
@@
$\mathbb{T}_{\alpha,\beta}:=\mathbb{I}(\kappa_{\alpha}^{*},\kappa_{\alpha+1})*\dot{\mathbb{T}}_{\alpha+1/2,\beta}$
@@
$\mathbb{T}(\kappa_{\alpha}^{*},\kappa_{\alpha+1})\times\prod_{\gamma\in(\alpha,\beta)}\mathbb{T}((\mathbb{P}_{\alpha,\gamma}^{G_{\alpha+1/2}},\dot{\leq}_{\alpha,\gamma}^{G_{\alpha+1/2}}),(\dot{\mathbb{Q}}_{\gamma},\dot{\leq}_{\gamma,0}))$
@@
We now prove (3). We will show that $\mathbb{I}(\kappa_{\alpha}^{*},\kappa_{\alpha+1})$ forces that $\dot{\mathbb{T}}_{\alpha+1/2,\beta}$ forces $\mathsf{LIP}(\check{\kappa}_{\alpha}^{*},(\check{\kappa}_{\alpha}^{*})^{+},\check{\kappa}_{\gamma})$ for every $\gamma\geq\alpha+1$, which clearly suffices. So let $H_{\alpha}$ be $\mathbb{I}(\kappa_{\alpha}^{*},\kappa_{\alpha+1})$-generic over $V[G_{\alpha}]$. In $V[G_{\alpha}*H_{\alpha}]$, the supercompactness of $\kappa_{\alpha+1}$ is preserved by any strongly $\kappa_{\alpha+1}-$directed closed forcing. In particular, it is preserved by $\mathbb{T}:=\prod_{\gamma\in(\alpha,\beta)}\mathbb{T}((\mathbb{P}_{\alpha,\gamma}^{G_{\alpha}*H_{\alpha}},\dot{\leq}_{\alpha,\gamma}^{G_{\alpha}*H_{\alpha}}),(\dot{\mathbb{Q}}_{\gamma},\dot{\leq}_{\gamma,0}))$. That forcing also preserves the fact that $\mathbb{T}(\kappa_{\alpha}^{*},\kappa_{\alpha+1})$ is a $\mathsf{LIP}$-forcing (which it is by Lemma 5.5 (5)). Ergo, $\mathbb{T}$ followed by $\mathbb{T}(\kappa_{\alpha}^{*},\kappa_{\alpha+1})$ forces $\mathsf{LIP}(\kappa_{\alpha}^{*},(\kappa_{\alpha}^{*})^{+},\kappa_{\gamma})$ for every $\gamma\geq\alpha+1$, since it collapses $\kappa_{\alpha+1}$ to become $(\kappa_{\alpha}^{*})^{+}$. The preservation of each $\kappa_{\gamma}$ follows just as in Lemma 6.5. ∎
@@
By Lemma 6.5, it suffices to show that whenever $\gamma$ is an ordinal with countable cofinality and $\alpha<\gamma$, then after forcing with $\mathbb{P}$ there are stationarily many $\beta\in\kappa_{\gamma}^{+}$ of cofinality $\kappa_{\alpha}^{*}$ which are not approachable (since the class of regular cardinals consists precisely of the $\kappa_{\alpha}^{*}$). So fix, in the ground model, a normal coloring $d\colon[\kappa_{\gamma}^{+}]^{2}\to\omega$ (it suffices to consider only one specific coloring by Fact 1.4)
@@
$\mathbb{P}_{\gamma}\cong\mathbb{P}_{\alpha}*\dot{\mathbb{I}}(\check{\kappa}_{\alpha}^{*},\check{\kappa}_{\alpha+1})*\dot{\mathbb{A}}(\check{\kappa}_{\alpha}^{*},\check{\kappa}_{\alpha+1})*\dot{\mathbb{P}}_{\alpha+1,\gamma}$

### v2

@@ 6. The Main Iteration
In this section we will prove the main theorem. Let us first explain the idea behind the construction: If $\kappa$ is supercompact, the poset $\mathbb{A}(\mu,\kappa)$ prepares the model as follows: Any further poset which forces $\mathsf{LIP}(\check{\mu},\check{\mu}^{+},(\check{\delta}_{n})_{n\in\omega})$ for some increasing sequence $(\delta_{n})_{n\in\omega}$ of regular cardinals also forces that there are stationarily many points in $(\sup_{n}\delta_{n})^{+}$ of cofinality $\mu$ which are not approachable. Given this poset, it makes sense to iterate the poset $\mathbb{A}$ for an increasing sequence of supercompact cardinals. The main problem, of course, is showing that enough instances of $\mathsf{LIP}$ are forced and the stationarity of the set of non-approachable points of cofinality $\mu$ is preserved. The solution to this problem consists of the following two ideas:
@@
    By Theorem 5.3, after forcing with $\mathbb{A}(\mu,\kappa)$ the stationarity of the set of non-approachable points of cofinality $\mu$ in $\delta^{*}:=(\sup_{n}\delta_{n})^{+}$ is witnessed by models $M$ such that there are totally $(M,\mathbb{P})$-generic conditions whenever $\mathbb{P}$ is a $\kappa-$directed closed forcing which preserves $\delta^{*}$, every $\delta_{n}$ and forces $\mathsf{LIP}(\check{\mu},\check{\mu}^{+},\check{\delta}_{n})$ for every $n\in\omega$.
@@
So the main obstacle is showing that $(\mathbb{P},\leq_{0})$ forces many instances of $\mathsf{LIP}$. However, in general working with the direct extension ordering on a Magidor iteration is complicated, so we work with the product of the termspace forcings instead, which suffices by the projection. To show that this forces $\mathsf{LIP}$, we need that our cardinal $\kappa$ is indestructably supercompact. By work of Hamkins and Shelah (see [HS98]), the indestructibility of a supercompact cardinal is destroyed by any nontrivial small forcing. Because of this, we follow every $\mathbb{A}$-iterand with the poset making the next supercompact cardinal indestructible. The following lemma easily follows from well-known work of Laver (see [Lav78]) together with our iteration lemma for strong directed closure (see Lemma 1.25):
@@
We will verify in a series of lemmas that $\mathbb{P}$ is as desired. Since $\mathbb{P}$ is not a set, but a class, “forcing with $\mathbb{P}$” is ill-defined – consider a forcing adding $\operatorname{On}$-many Cohen reals. However, in our case we will show the following:
@@
In particular, for any $\alpha<\beta$, forcing with $\dot{\mathbb{P}}_{\alpha,\beta}^{G}$ (where $G$ is $\mathbb{P}_{\alpha}$-generic) does not add any new bounded subsets of $\kappa_{\alpha}$ or $\kappa_{\alpha}^{+}$. We may therefore construct a forcing extension “by $\mathbb{P}$” as follows: Let $G\subseteq\mathbb{P}$ be a filter which intersects every set-sized dense subclass of $\mathbb{P}$. For any $\alpha\in\operatorname{On}$, $G\upharpoonright\alpha:=\{p\upharpoonright\alpha\;|\;p\in G\}$ is a generic filter over $\mathbb{P}_{\alpha}$. Let $H^{*}(\kappa_{\alpha})$ be equal to $H(\kappa_{\alpha})$ as computed in $V[G_{\alpha}]$. It follows that, for any $\beta>\alpha$, $H^{*}(\kappa_{\alpha})$ is equal to $H(\kappa_{\alpha})$ as computed in $V[G_{\beta}]$, since $\dot{\mathbb{P}}_{\alpha,\beta}^{G_{\alpha}}$ does not add any new elements of $H(\kappa_{\alpha})$. Lastly, let
$V[G]:=\bigcup_{\alpha\in\operatorname{On}}H^{*}(\kappa_{\alpha}).$
It follows that $V[G]$ is a model of $\mathsf{ZFC}$, since any axiom can be verified locally. For example, the powerset axiom holds in $V[G]$ as follows: Any $x\in V[G]$ is in some $H^{*}(\kappa_{\alpha})$. Thus, any subset of $x$ is an element of $H^{*}(\kappa_{\alpha})$ as well. Therefore, the powerset of $x$ in $V[G]$ is equal to the powerset of $x$ in $V[G\upharpoonright\alpha]\subseteq V[G\upharpoonright\alpha+1]$ and thus a member of $H^{*}(\kappa_{\alpha+1})\subseteq V[G]$.
Moreover, for any $\alpha\in\operatorname{On}$, $H(\kappa_{\alpha})$ as computed in $V[G]$ is equal to $H^{*}(\kappa_{\alpha})$ and thus equal to $H(\kappa_{\alpha})$ as computed in $V[G\upharpoonright\alpha]$. Therefore, any bounded assertion can be verified by proving that it holds in $V[G\upharpoonright\beta]$ for all sufficiently large $\beta$.
For more information on class-sized forcing extensions, see e.g. the proof of Easton’s Theorem in Jech’s textbook[Jec03, Page 235]. Class forcing can be avoided at the cost of increasing the large cardinal assumptions by instead starting with a sequence $(\kappa_{\alpha})_{\alpha<\delta}$, where $\delta$ is inaccessible and then performing the set-sized iteration up to $\delta$, whence the conclusion of the main theorem holds in $V_{\delta}$.
@@
So we have shown that all $\kappa_{\alpha}$ and $\kappa_{\alpha}^{*}$ remain cardinals. Clearly, whenever $\mu$ does not have this form, $\mu$ is collapsed. So the above class contains precisely all cardinals in the forcing extension. ∎
@@
We want to show that $\mathbb{P}$ gives us exactly the model we need. The proof idea is as follows: The regular cardinals of the forcing extension are precisely the $\kappa_{\alpha}$ for $\alpha$ a successor and $\kappa_{\alpha}^{+}$ for $\alpha$ a limit. For any such cardinal, assuming enough instances of $\mathsf{LIP}$, $\mathbb{A}(\kappa_{\alpha}^{*},\kappa_{\alpha+1})$ forces the existence of stationarily many non-approachable points of cofinality $\kappa_{\alpha}^{*}$ in any successor of a singular cardinal $\kappa_{\alpha+1}$ of countable cofinality. Then all that is left is to show that the tail of the iteration does not destroy that stationarity (since it cannot make points of cofinality $\kappa_{\alpha}^{*}$ approachable due to its closure). For this we use that we have used a guessing function and thus added many partially generic filters to many strongly $\kappa_{\alpha}-$directed closed forcings, namely, those forcings which force enough instances of $\mathsf{LIP}$. We now show that the direct extension ordering on the tail of the iteration can be projected onto from one of those suitable forcings.
@@
$\mathbb{T}_{\alpha,\beta}:=\mathbb{I}(\kappa_{\alpha}^{*},\kappa_{\alpha+1})*\dot{\mathbb{T}}_{\alpha+1/2,\beta}.$
@@
$\mathbb{T}(\kappa_{\alpha}^{*},\kappa_{\alpha+1})\times\prod_{\gamma\in(\alpha,\beta)}\mathbb{T}((\mathbb{P}_{\alpha,\gamma}^{G_{\alpha+1/2}},\dot{\leq}_{\alpha,\gamma}^{G_{\alpha+1/2}}),(\dot{\mathbb{Q}}_{\gamma},\dot{\leq}_{\gamma,0})).$
@@
We now prove (3). We will show that $\mathbb{I}(\kappa_{\alpha}^{*},\kappa_{\alpha+1})$ forces that $\dot{\mathbb{T}}_{\alpha+1/2,\beta}$ forces $\mathsf{LIP}(\check{\kappa}_{\alpha}^{*},(\check{\kappa}_{\alpha}^{*})^{+},\check{\kappa}_{\gamma})$ for every $\gamma\geq\alpha+1$, which clearly suffices. So let $H_{\alpha}$ be $\mathbb{I}(\kappa_{\alpha}^{*},\kappa_{\alpha+1})$-generic over $V[G_{\alpha}]$. In $V[G_{\alpha}*H_{\alpha}]$, the supercompactness of $\kappa_{\alpha+1}$ is preserved by any strongly $\kappa_{\alpha+1}-$directed closed forcing. In particular, it is preserved by $\mathbb{T}:=\prod_{\gamma\in(\alpha,\beta)}\mathbb{T}((\mathbb{P}_{\alpha,\gamma}^{G_{\alpha}*H_{\alpha}},\dot{\leq}_{\alpha,\gamma}^{G_{\alpha}*H_{\alpha}}),(\dot{\mathbb{Q}}_{\gamma},\dot{\leq}_{\gamma,0}))$. By Lemma 5.6 (5), $\mathbb{T}(\kappa_{\alpha}^{*},\kappa_{\alpha+1})$ is a $\mathsf{LIP}$-forcing, which is preserved by $\mathbb{T}$. Ergo, $\mathbb{T}$ followed by $\mathbb{T}(\kappa_{\alpha}^{*},\kappa_{\alpha+1})$ forces $\mathsf{LIP}(\kappa_{\alpha}^{*},(\kappa_{\alpha}^{*})^{+},\kappa_{\gamma})$ for every $\gamma\geq\alpha+1$, since it collapses $\kappa_{\alpha+1}$ to become $(\kappa_{\alpha}^{*})^{+}$. The preservation of each $\kappa_{\gamma}$ follows just as in Lemma 6.5. ∎
@@
By Lemma 6.5, it suffices to show that whenever $\gamma$ is an ordinal with countable cofinality and $\alpha<\gamma$, then after forcing with $\mathbb{P}$ there are stationarily many $\beta\in\kappa_{\gamma}^{+}$ of cofinality $\kappa_{\alpha}^{*}$ which are not approachable, since the class of regular cardinals consists precisely of the $\kappa_{\alpha}^{*}$. So fix, in the ground model, a normal coloring $d\colon[\kappa_{\gamma}^{+}]^{2}\to\omega$. It suffices to consider only one specific coloring by Fact 1.5.
@@
$\mathbb{P}_{\gamma}\cong\mathbb{P}_{\alpha}*\dot{\mathbb{I}}(\check{\kappa}_{\alpha}^{*},\check{\kappa}_{\alpha+1})*\dot{\mathbb{A}}(\check{\kappa}_{\alpha}^{*},\check{\kappa}_{\alpha+1})*\dot{\mathbb{P}}_{\alpha+1,\gamma}.$